OpenStudy (anonymous):

CAN SOMEONE PLEASE SOLVE THIS IN AS MUCH DETAIL AS POSSIBLE? x^2 + x - 12 = 0

OpenStudy (akshay_budhkar):

x^2+4x-3x-12=0 (splitting x) x(x+4)-3(x+4)=0 taking common x and -3 respectively (x-3)(x+4)=0 taking (x-4) common a*b=0 then a=0 or b=0 or both so we have x+4=0 or x-3=0 therefore x=-4 or x=3

OpenStudy (akshay_budhkar):

in step 3 make it x+4 common.. typo

OpenStudy (anonymous):

thank you! :D